OpenStudy (anonymous):

describe how to transform (^3 sqrt x^4)^5 into an expression with a rational exponent. make sure you respond with complete sentences

OpenStudy (paounn):

you know that in general, \[\sqrt[n]{x^m} = x^{m/n}\] Your expression becomes \[(\sqrt[3]{x^4})^5 = (x^{3/4})^5 = x^{(3*5)/4} = x^ {15/4} \] Notice the use of brakets. You should also remember that \[(a^n)^m = a^{n*m}\]

OpenStudy (anonymous):

so thats it ?

OpenStudy (anonymous):

that cant be the whole equation if so thanks so much for your help